OpenStudy (anonymous):

Find the distance between the two points. (-6, 1) (3, 1) a. 7 b. 10 c. 8 d. 9

OpenStudy (anonymous):

d

OpenStudy (anonymous):

d

OpenStudy (anonymous):

(-6-3)^2+(1-1)^2=9^2 sqrt9^2=9